OpenStudy (anonymous):

find a constant c such that the graph of x^2+11x+c in the xy plane has its vertex on the line y=x the hint said "complete the square"

OpenStudy (anonymous):

Let's see... try adding 30 and subtracting 30, then getting (x+5)(x+6)+c-30... well, that's completing the square. What I'd do is find y'=2x+11=0, x=-11/2, and then (-11/2)^2+11(-11/2)+c=-11/2. From here solving for c should be trivial.

OpenStudy (anonymous):

x^2+11x+121/4-121/4+c (x+11/2)^2+c-121/4 c-121/4=0 c=121/4

OpenStudy (anonymous):

or x^2+11x+c=x x^2+10x+c=0 ax^2+bx+c=0 delta=b^2-4ac delta=0 y=x line tangent of parabola it is already said in the question 10^2-4*1*c=0 c=25